OpenStudy (anonymous):

use natural logarithms to solve: e^x=18 please show work

OpenStudy (mertsj):

Take the natural log of both sides.

OpenStudy (anonymous):

ln e^x = ln18 x = ln18 x approxamately equals 2.89

OpenStudy (mertsj):

\[e^x=18\] \[\ln e^x=\ln 18\] \[x \ln e=\ln 18\]

OpenStudy (mertsj):

Since lne = 1, x = ln 18

OpenStudy (anonymous):

okay thanks guys!